Use the this set of data and table to answer the following questions

Grade Count
A 10
B 2
C 1
D 0
F 2

What is the probability that a student received grades A AND B?

If a committee with 2 student members is to be formed, what is the probability of forming a committee with one A grade student and one F grade student?

If a committee with 3 student members is to be formed, what is the probability of forming a committee with two A grade students and one B grade student?
Answer #1

total students = 10+2+1+0+2 = 15

a) probability that a student received grades A AND B =10/15*2/15=0.0889

b) the probability of forming a committee with one A grade student and one F grade student = 10C1*2C1/15C2 = 10*2/105=0.1905

c)  probability of forming a committee with two A grade students and one B grade student=10C2*2C1/15C3 = 0.19780